# Diffusion Models for Tabular Data Imputation and Synthetic Data Generation

### v1

@@ 4.2. Model Overview
As shown in Figs. 1, 2 and 3, the denoising model is based on an encoder-decoder structure. The encoder obtains a representation of each feature in two steps. Initially, a columnar embedding projects individually all the heterogeneous features (continuous or categorical) in the same homogeneous and dense latent space. Then, a conditional encoder-decoder transformer embedding enhances the latent representation of features by considering their interactions with each other and with the condition. For the columnar embedding of categorical and numerical features, we utilize an embedding layer and a linear-ReLU activation layer, respectively. The type of task (regression or classification) dictates the choice of embedding for the target (y).
The conditional encoder-decoder transformer (detailed in Fig. 2) incorporates the output of columnar embedding through specialized masking. More specifically, the encoder focuses on the variables $\bar{x}^{M}$ and $y$ (i.e. conditioning variables), while the decoder attends to all the available information, $x^{M}$, $\bar{x}^{M}$ and $y$. In this setup, the encoder output provides context from conditioning variables not involved in the forward diffusion process, allowing the decoder to derive representations for those that are. Features involved in the forward diffusion process are denoted with $\texttt{Mask}=1$, and those that aren’t with $\texttt{Mask}=0$. Consequently, the encoder takes in $1-\texttt{Mask}$. When the original dataset contains missing values, a new MissingMask is introduced. More concretely, the MissingMask is combined with the conditioning Mask to remove any missing value from the condition. This mechanism allows the model to work without any type of missing data preprocessing.
@@
We note that the encoder’s output is converted into a set of attention vectors. Specifically, the decoder integrates these vectors into its attention mechanism (Fig. 3) as the K and V vectors described in [54]. This allows the decoder to focus on the relevant features of the condition and their interactions with the masked features. This conditioning mechanism is more general and exhibits less learning bias compared to recent approaches in the literature on diffusion models for tabular data, such as TabDDPM and TabCSDI, where the condition embedding is only added to the masked feature embedding. This advantage holds true even in scenarios where the encoder processes only one variable (the target variable of a supervised task). Here, the transformer encoder functions similarly to an MLP, but its output continues to be used in the decoder’s attention mechanism, enhancing overall performance (see discussion on ablation study in Sec. 6).
Figure 2. Conditional Transformer Encoder: In the input of the encoder-decoder transformer, the embedding of a conditioning feature is represented by a fully filled square ($\texttt{Mask}=0$), and the embedding of a masked feature is depicted by a color-dashed square ($\texttt{Mask}=1$). The encoder’s task is to learn a representation of the conditioning features, using only those marked with $\texttt{Mask}=0$. On the other hand, the decoder focus on improving the representation of the masked features ($\texttt{Mask}=1$), utilizing all available features, both conditioning and masked, along with the context generated by the encoder. The square filled with colored rhombuses in the output of the decoder correspond with the improved representation of the masked features (i.e., $\texttt{Mask}=1$). When the original dataset contains missing values, a new MissingMask is introduced. The MissingMask is combined with the conditioning Mask to remove missing value from the condition.
@@
Figure 3. Conditioning Attention Mechanism: The condition embedding produced by the transformer encoder is used in the decoder attention mechanism. More in details, the condition embedding plays the roles of K and V whereas the feature embedding plays the role of Q.
@@
The final latent feature representation is obtained by summing the conditional encoder-decoder transformer output with the timestep embedding, which is derived by projecting the sinusoidal time embedding [38, 10] into the transformer embedding dimension, using a linear layer followed by the Mish activation function [35]. Last, this representation is decoded to produce the output. Each feature has its own decoder consisting of two successive blocks integrating a linear transformation, normalization, and ReLU activation. Depending on the feature type (numerical or categorical), an additional linear layer is appended with either a singular output for numerical features or multiple outputs, corresponding to the number of classes for categorical ones.
The model is trained by minimizing a sum of mean-squared error $L^{simple}_{t}$ (Eq. 8) for the Gaussian diffusion term and the KL divergences $L^{i}_{t}$ for each multinomial diffusion term (Eq. 1).
$L^{MTabGen}_{t}=$ $\dfrac{L^{simple}_{t}(\texttt{Mask})}{K_{num}}+\dfrac{\sum_{i\leq K_{cat}}L^{i}_{t}(\texttt{Mask})/Cl_{i}}{K_{cat}}$
@@
where $L^{simple}_{t}(\texttt{Mask})$ and $L^{i}_{t}(\texttt{Mask})$ means that the loss functions are computed taking into account only the prediction error on variables affected by the forward diffusion process (i.e. $x^{M}$), $Cl_{i}$ is the number of classes of the $i$-th categorical variable.
@@ 4.3. Dynamic Conditioning
A key feature of the proposed solution is that the split between $x^{M}$ and $\bar{x}^{M}$ does not have to be fixed for all the row $i$ in the dataset. The transformer encoder can manage mask with an arbitrary number of zeros/ones, so we can dynamically alter the split between $x^{M}$ and $\bar{x}^{M}$ by just producing a new mask. In the extreme scenario, we can generate a new mask $\texttt{Mask}_{i}$ for each row $i$. During training, the number of ones in $\texttt{Mask}_{i}$ (i.e. the number of features to be included in the forward diffusion process) is uniformly sampled from the interval $[1,K_{num}+K_{cat}]$. A model that has been trained in this manner can then be used for both tasks, that is, generation of synthetic data ($\texttt{Mask}_{i}=1$ for all the $K_{num}+K_{cat}$ features and for any dataset index $i$) and imputation of missing values (for each $i$, $\texttt{Mask}_{i}=1$ for the feature to impute). As discussed, when the original dataset contains missing values, a new MissingMask is introduced and combined with the conditioning Mask to remove any missing value from the condition. The MissingMask is fixed and constant during the training phase. This setup allows for more flexible conditioning scenarios.
@@
  • When $\bar{x}^{M}=\emptyset$, $p\left(x^{M}_{t-1}\middle|x^{M}_{t},\bar{x}^{M},y\right)=p\left(x_{t-1}\middle|x_{t},y\right)$ our model aligns with TabDDPM, generating synthetic data influenced by the target distribution.
@@ 5.1. Data
  • Forest Cover Type [7]: In this multi-class classification task with seven categories, cartographic variables are used to predict the forest cover type. The first eight features of the dataset are continuous, whereas the last two are categorical with four and 40 levels respectively.
@@ 5.2. Baselines
  • TVAE555 We use the implementation provided by https://sdv.dev/ [59]: A variational autoencoder adapted for mixed-type tabular data.
  • CTGAN5 [59]: A conditional GAN for synthetic tabular data generator.
@@
  • missForest666 We use the implementation provided by https://github.com/vanderschaarlab/hyperimpute [48]: Iterative method based on random forests to predict and fill in missing values.
  • GAIN6 [60]: GAN model for tabular missing value imputation.
  • HyperImpute6 [23]: Iterative imputation algorithm using both regression and classification methods based on linear models, trees, XGBoost, CatBoost and neural nets.
  • Miracle6 [31]: Missing imputation algorithm using a causal deep learning approach.
@@ 5.3.1. Machine Learning efficiency
The Machine Learning efficiency measures the performance degradation of classification or regression models trained on synthetic data, and then tested on real data. The basic idea is to use a ML discriminative model to evaluate the utility of synthetic data provided by a generative model. As demonstrated by Kotelnikov et al. [30], a strong ML model allows to obtain more stable and consistent conclusions on the performances of the generative model. Based on this intuition, we consider 4 different ML models: XGBoost [8], CatBoost [42], LightGBM [27] and MLP. We introduce an initial fine-tuning step, during which we derive the best hyperparameter configuration using the Optuna library [2]. Specifically, we perform 100 iterations to fine-tune the model’s (XGBoost, CatBoost, LightGBM and MLP) hyperparameters on each dataset’s real data within the benchmark. Every hyperparameter configuration for ML model is cross-validated, using a five-fold split. The complete hyperparameter search space is shown in appendix A Table 7.

### v2

@@ 4.2. Model Overview
Figure 1. Denoising function $dn_{\theta}$. The function takes noisy data representations $x_{t}^{n}$, $x_{t}^{c}$, conditioning values $x^{\text{cond}}=x_{0}^{n},x_{0}^{c}$ and conditioning target $y_{0}$, a conditional mask Mask and current timestep $t$ as inputs. The Columnar Embedding layer projects numerical and categorical features into a shared latent space, while the Encoder-Decoder Transformer (detailed in Fig. 2) refines this representation by modeling intra- and inter-feature relationships. Outputs are generated through a Columnar Decoder, predicting the estimated noise for numerical features ($\hat{\epsilon}^{n}_{t}$) and logits for categorical features ($\hat{x}^{c}_{0}$).
MTabGen utilizes an encoder-decoder transformer architecture as the denoising function to refine the representation of noisy data and produce clean samples through reverse diffusion. The workflow is depicted in Figure 1, with the following key components: the columnar embedding, the encoder-decoder transformer, and the output decoders for each feature type.
4.2.1. Denoising Function Architecture
The denoising function $dn_{\theta}$ is responsible for reconstructing the original data representation from noisy samples. Figures 1, 2, and 3 illustrate its detailed workflow and components. The function follows an encoder-decoder model structure and takes as inputs the current noisy representations of numerical and categorical data ($x^{n}_{t}$ and $x^{c}_{t}$) at timestep $t$, a conditional mask (Mask) identifying features involved in the forward diffusion process, the conditioning values ($x^{\text{cond}}$), the target ($y$), and the current timestep ($t$).
The primary objective of the denoising function is to produce two outputs at each time step: the estimated noise ($\hat{\epsilon}^{n}_{t}$) for numerical features and the logits ($\hat{x}^{c}_{0}$) for categorical features, representing their predicted original distribution. These estimated outputs are then used by Eqs. 7 and 10, respectively, to perform the reverse denoising at each timestep $t$ as described in the following sections.
4.2.2. Columnar Embedding
To facilitate the learning of the denoising function within the transformer model, a columnar embedding projects both numerical and categorical features of $x^{\text{mask}}$ and $x^{\text{cond}}$ into a shared latent space. In the case of categorical features, an embedding layer maps each categorical value into a dense vector representation, whereas numerical features are embedded using a linear transformation followed by a ReLU activation function. The embedding of the target variable $y$ depends on the supervised task. For regression, $y$ is treated as a continuous variable, while for classification, it is processed through an embedding layer similar to other categorical features.
4.2.3. Conditional Encoder-Decoder Transformer
Figure 2. Conditional Transformer Encoder-Decoder model. The encoder takes three inputs: noisy data embedding, condition embedding, and a conditional mask (Mask). The embedding of a conditioning feature is represented by a fully filled square, while the embedding of a noisy feature is depicted as a dashed-colored square. The effective mask (EffMask) combines missing and conditional masks, allowing the encoder to process conditioning features and target that are unaffected by forward diffusion and without missing values, learning a conditioning context vector for the decoder. The decoder then refines the representation of masked features using all available information, including the encoder context, both masked and conditioning features and targets. The final representation of the masked features is depicted as squares filled with colored rhombuses.
The workflow, illustrated in Fig. 2, consists of three inputs, two intermediate stages—masking and transformer—and a final output, arranged from bottom to top. The three inputs are: (1) the columnar embedding of the noisy data, referred to as the noisy embedding, (2) the columnar embedding of the conditioning data, termed the condition embedding, and (3) the conditional mask, denoted as Mask . In the conditional mask, features involved in the forward diffusion process are indicated by $\texttt{Mask}=1$, while features not involved are marked as $\texttt{Mask}=0$. If the original dataset contains missing values, an additional mask, termed MissingMask, is introduced, where $\texttt{MissingMask}=1$ for missing values, as shown in Fig. 2.
[…]
In the transformer phase, the encoder exclusively processes the conditioning features characterized by $\texttt{EffMask}=0$, which correspond to features unaffected by the forward diffusion process and those that do not contain missing values. Simultaneously, the decoder processes two distinct types of inputs: (1) a combination of noisy data, which includes features involved in the forward diffusion process or containing missing values ($\texttt{EffMask}=1$), and conditioning features ($\texttt{EffMask}=0$), and (2) the output from the encoder . This setup enables the encoder to provide contextual information from the conditioning features, allowing the decoder to generate refined representations of the features that require denoising. The output of the decoder corresponds to the final output stage, as shown in Fig. 2.
Decoder and Attention Mechanism
The decoder attends to all conditioning features, using the context provided by the encoder. The attention mechanism, depicted in Figure 3, incorporates the output of the encoder as the key ($\bf{K}$) and value ($\bf{V}$) vectors, while the features, including the noisy data, serves as the query ($\bf{Q}$) vector, as described in Vaswani et al. [59]. This enables the decoder to refine masked feature representations by selectively focusing on interactions between conditioning variables and noisy data.
@@ 4.2.3. Conditional Encoder-Decoder Transformer
This attention-based conditioning mechanism is more general and exhibits less learning bias compared to recent approaches in the literature on diffusion models for tabular data, such as TabDDPM and TabCSDI, where the condition embedding is only added to the masked feature embedding. This advantage holds true even in scenarios where the encoder processes only one variable (the target variable of a supervised task). Here, the transformer encoder functions similarly to an MLP, but its output continues to be used in the decoder’s attention mechanism, enhancing overall performance (see discussion on ablation study in Sec. 6).
@@
Figure 3. Conditioning Attention Mechanism: The condition embedding produced by the transformer encoder is used in the decoder attention mechanism. More in details, the condition embedding plays the roles of $\bf{K}$ and $\bf{V}$ whereas the feature embedding plays the role of $\bf{Q}$.
4.2.4. Output Denoising Parameters
@@
Following with the description of the workflow of Fig. 1, the final latent representation of the features to denoise is obtained by summing the conditional encoder-decoder transformer output with the timestep embedding, which is derived by projecting the sinusoidal temporal embedding [41, 11] into the transformer embedding dimension, using a linear layer followed by the Mish activation function [37]. Last, this representation is decoded to produce the output. Each feature has its own decoder consisting of two successive blocks integrating a linear transformation, normalization, and ReLU activation. Depending on the feature type (numerical or categorical), an additional linear layer is appended with either a singular output for numerical features or multiple outputs, corresponding to the number of classes for categorical ones.
4.3. Sampling or Reverse Diffusion Process
The primary goal of the sampling or reverse diffusion process is to convert a set of samples from either Gaussian (for numerical variables) or uniform distributions (for categorical variables) into clean samples that represent a complex real-world distribution. The sampling proceeds iteratively, moving from timestep $t=T$, where the distribution is simple and noisy, to timestep $t=0$, where the representation becomes clean and realistic.
Algorithm 1 Sampling or Reverse diffusion process
1:Inputs: Denoising function $dn_{\theta}$, Conditional Mask Mask, Conditioning Value $x^{\text{cond}}$, $y$ and number of Timesteps $T$
2:Preprocess the conditioning value $x^{\text{cond}}$, $y$.
3:Initialize noise:
4:     Sample initial numerical gaussian noise: $x^{n}_{T}\sim\mathcal{N}\left(x^{n}_{T};0,I\right)$.
5:     Sample initial categorical uniform noise: $x^{c}_{T}\sim Cat\left(x^{c}_{T};1/Cl\right)$.
6:for t=T, T-1, …, 2, 1, 0 do
7:     $\hat{\epsilon}^{n}_{t},\hat{x}^{c}_{0}=dn_{\theta}\left(x^{n}_{t},\,x^{c}_{t},\,\texttt{Mask},\,x^{\text{cond}},\,y,\,t\right)$.
8:     Perform a reverse Gaussian diffusion step using Eq. 7 to compute $x^{n}_{t-1}$ with $\hat{\epsilon}^{n}_{t}$ and $x^{n}_{t}$.
9:     Perform a reverse multinomial diffusion step using Eq. 10 to compute $x^{c}_{t-1}$ with $\hat{x}^{c}_{0}$ and $x^{c}_{t}$.
10:end for
11:Postprocess $x^{n}_{0}$ and $x^{c}_{0}$ to reverse the preprocessing from state 2.
The iterative reverse diffusion process for the generation of one sample is described in Algorithm 1. This algorithm requires four inputs: the denoising function, the conditional mask, the conditioning values, and the number of timesteps. The denoising function, implemented as a neural network, removes noise from the data and is shared across all variables to enhance its capacity to learn complex relationships. The conditional mask distinguishes the conditioning variables, $x^{\text{cond}}$ and $y$, from those involved in the forward diffusion process that require denoising. Variables requiring denoising are marked with $\texttt{Mask}=1$, while those that do not are marked with $\texttt{Mask}=0$. The conditioning values are given by $x^{\text{cond}}$ and $y$ at time $t=0$. The number of timesteps dictates how many reverse diffusion steps must be performed.
First, a consistent preprocessing procedure is applied to the conditioning values, using the Gaussian quantile transformation from the scikit-learn library [44] for numerical features and ordinal encoding for categorical ones. Then, at time $t=T$, the process initiates by sampling the initial noise for each dataset feature separately: Gaussian noise for numerical features and uniform noise for categorical features. The iterative denoising process then proceeds from $t=T$ to $t=0$. Each iteration involves three key actions: first, the denoising function $dn_{\theta}$ estimates $\hat{\epsilon}^{n}_{t}$ for continuous features and $\hat{x}^{c}_{0}$ for categorical features. Next, a reverse Gaussian step and a reverse multinomial step are performed using Eqs. 7 and 10 to estimate $x^{n}_{t-1}$ and $x^{c}_{t-1}$, respectively. These actions are repeated iteratively until the final outputs $x^{n}_{0}$ and $x^{c}_{0}$ are reached. Finally, a postprocessing step is applied to the final outputs $x^{n}_{0}$ and $x^{c}_{0}$ to reverse the initial preprocessing. It is important to note that the preprocessing of conditioning values and postprocessing of outputs enhances the learning rate without impacting the quality of the final results. We tested various normalization methods for numerical features, such as Standard Scaler and Min-Max Scaler from scikit-learn, and found the results to be statistically equivalent.
4.4. Training process
The training process for the denoising function is outlined in Alg. 2. The primary aim is to train the neural network $dn_{\theta}$ used in the state 7 of Alg. 1. This network is designed to leverage noisy representations of numerical and categorical features $x^{n}_{t},x^{c}_{t}$, along with the conditional mask Mask, conditioning values, and current timestep $t$, to estimate both $\hat{\epsilon}^{n}_{t}$ and $\hat{x}^{c}_{0}$. For numerical features, $\hat{\epsilon}^{n}_{t}$ represents the estimated noise introduced during the forward diffusion process between timesteps $t-1$ and $t$. For categorical features, $\hat{x}^{c}_{0}$ serves as an estimate of the original categorical feature a $t=0$.
Algorithm 2 Training process
1:Inputs: Denoising function $dn_{\theta}$, number of Timesteps $T$, original data $x^{n}_{0},x^{c}_{0}$
2:Initialize the weights $\theta$ of the Denoising function $dn_{\theta}$.
3:Preprocess the original data $x^{n}_{0},x^{c}_{0}$
4:for iteration=1, 2, …, N do
5:     Sample a batch of original features, $x^{n}_{0},\,\,x^{c}_{0}$.
6:     Sample a batch of timesteps $t$ uniformly distributed within the interval $\left[0,T\right]$.
7:     for each numerical features $x^{n}_{0}$ do
8:         Use the forward Gaussian diffusion process of Eq. 5 (with reparametrization trick) to obtain $x^{j_{n}}_{t}$ and $\epsilon^{n}_{t}$.
9:     end for
10:     for each categorical feature $x^{c}_{0}$ do
11:         Use the forward multinomial diffusion process of Eq. 9 to obtain $x^{c}_{t}$.
12:     end for
13:     Generate the new conditional Mask
14:     Compute the denoising function outputs: $\hat{\epsilon}^{n}_{t},\hat{x}^{c}_{0}=dn_{\theta}\left(x^{n}_{t},x^{c}_{t},\texttt{Mask},x^{n}_{0},x^{c}_{0},t\right)$.
15:     Use Eq. 8 to compute $L_{simple}\left(\texttt{Mask}\right)$ for numerical features with $\texttt{Mask}=1$; using $\epsilon^{n}_{t}$ and $\hat{\epsilon}^{n}_{t}$
16:     Use Eq. 1 to compute $L_{t}\left(\texttt{Mask}\right)$ for categorical features with $\texttt{Mask}=1$; using $x^{c}_{0}$, $\hat{x}^{c}_{0}$ and $x^{c}_{t}$
17:     Compute the total loss function of Eq. 11.
18:     Minimize the total loss function and update the denoising function weights.
19:end for
The first steps involve initializing the model weights and preprocessing the original data as discussed in the previous section, followed by an iterative learning process. In each iteration, a batch of original features $x^{n}_{0},\,\,x^{c}_{0}$ is sampled from the original dataset. For each element in the batch, a timestep $t$ is then sampled uniformly from the interval $\left[0,T\right]$. Using the forward Gaussian diffusion process, targets for numerical features are computed. Specifically, for each numerical feature, the reparametrization trick is applied to Eq. 5 to compute $x^{n}_{t}$, that is the noisy representation of the numerical feature at time $t$, and $\epsilon^{n}_{t}$, the corresponding noise. Similarly, the forward multinomial diffusion process of Eq. 9 is used to obtain $x^{c}_{t}$, the noisy representation of the categorical feature a time $t$. Next, a new mask is dynamically generated to select a subset of features that will serve as the conditioning input. With all the all the $dn_{\theta}$ inputs ($x^{n}_{t}$, $x^{c}_{t}$, Mask, $x^{n}_{0}$, $x^{c}_{0}$, $t$) and targets ($\epsilon^{n}_{t}$ and $x^{c}_{0}$) in place, the loss function is then calculated.
The model is trained by minimizing the following total loss function:
(11) $L^{MTabGen}_{t}=\dfrac{L^{simple}_{t}(\texttt{Mask})}{K_{num}}+\dfrac{\sum_{i\leq K_{cat}}L^{i}_{t}(\texttt{Mask})/Cl_{i}}{K_{cat}}$
The total loss function $L^{MTabGen}_{t}$ includes two main components:
  • Loss for numerical features: $L^{simple}_{t}$ (defined in Eq. 8) computes the mean squared error between the true noise $\epsilon^{n}_{t}$ and the predicted noise $\hat{\epsilon}^{n}_{t}$ introduced in the forward Gaussian diffusion process between the steps $t-1$ and $t$.
  • Loss for categorical features: $L^{i}_{t}$ (defined in Eq. 1) calculates the Kullback-Leibler divergence between the posterior of the forward process $q\left(x_{t-1}^{c}\middle|x_{t}^{c},x_{0}^{c}\right)$ and parametrized reverse diffusion process $\hat{p}\left(x_{t-1}^{c}\middle|x_{t}^{c}\right)$. Following the approach in [20], the $\hat{p}\left(x_{t-1}^{c}\middle|x_{t}^{c}\right)$ is approximated by $q\left(x_{t-1}^{c}\middle|x_{t}^{c},\hat{x}^{c}_{0}\right)$ using Eq. 10.
@@ 4.4. Training process
$L^{simple}_{t}(\texttt{Mask})$ and $L^{i}_{t}(\texttt{Mask})$ means that the loss functions are computed taking into account only the prediction error on variables affected by the forward diffusion process (i.e. $\texttt{Mask}=1$). Here, $Cl_{i}$ is the number of classes of the $i$-th categorical variable.
@@ 4.5. Dynamic Conditioning
A key feature of the proposed solution is that the split between $x^{\text{mask}}$ and $x^{\text{cond}}$ does not have to be fixed for every row $i$ in the dataset. The transformer encoder can manage mask with an arbitrary number of zeros/ones, so we can dynamically alter the split between $x^{\text{mask}}$ and $x^{\text{cond}}$ by just producing a new mask. In the extreme scenario, we can generate a new mask $\texttt{Mask}_{i}$ for each row $i$. During training, the number of ones in $\texttt{Mask}_{i}$ (i.e. the number of features to be included in the forward diffusion process) is uniformly sampled from the interval $[1,K_{num}+K_{cat}]$. A model that has been trained in this manner can then be used for both tasks, that is, generation of synthetic data ($\texttt{Mask}_{i}=1$ for all the $K_{num}+K_{cat}$ features and for any dataset index $i$) and imputation of missing values (for each $i$, $\texttt{Mask}_{i}=1$ for the feature to impute). As discussed, when the original dataset contains missing values, a new MissingMask is introduced and combined with the conditional Mask to remove any missing value from the condition. The MissingMask is fixed and constant during the training phase. This setup allows for more flexible conditioning scenarios.
@@
  • When $x^{\text{cond}}=\emptyset$, $p\left(x^{\text{mask}}_{t-1}\middle|x^{\text{mask}}_{t},x^{\text{cond}},y\right)=p\left(x_{t-1}\middle|x_{t},y\right)$ our model aligns with TabDDPM, generating synthetic data influenced by the target distribution.
@@ 5.1. Data
  • Forest Cover Type [7]: In this multi-class classification task with seven categories, cartographic variables are used to predict the forest cover type. The first eight features of the dataset are continuous, whereas the last two are categorical with 4 and 40 levels respectively.
@@ 5.2. Baselines
  • TVAE444 We use the implementation provided by https://sdv.dev/ [64]: A variational autoencoder adapted for mixed-type tabular data.
  • CTGAN4 [64]: A conditional GAN for synthetic tabular data generator.
@@
  • missForest555 We use the implementation provided by https://github.com/vanderschaarlab/hyperimpute [53]: Iterative method based on random forests to predict and fill in missing values.
  • GAIN5 [65]: GAN model for tabular missing value imputation.
  • HyperImpute5 [25]: Iterative imputation algorithm using both regression and classification methods based on linear models, trees, XGBoost, CatBoost and neural nets.
  • Miracle5 [33]: Missing imputation algorithm using a causal deep learning approach.
@@ 5.3.1. Machine Learning efficiency
The Machine Learning efficiency measures the performance degradation of classification or regression models trained on synthetic data, and then tested on real data. The basic idea is to use a ML discriminative model to evaluate the utility of synthetic data provided by a generative model. As demonstrated by Kotelnikov et al. [32], a strong ML model allows to obtain more stable and consistent conclusions on the performances of the generative model. Based on this intuition, we consider 4 different ML models: XGBoost [9], CatBoost [45], LightGBM [29] and MLP. We introduce an initial fine-tuning step, during which we derive the best hyperparameter configuration using Bayesian optimization and Optuna library [2]. Specifically, we perform 100 iterations to fine-tune the model’s (XGBoost, CatBoost, LightGBM and MLP) hyperparameters on each dataset’s real data within the benchmark. Every hyperparameter configuration for ML model is cross-validated, using a five-fold split. The complete hyperparameter search space is shown in Appendix A Table 9.